Rise and run. If the rise is
3/2 and the run is 5, then what is
the ratio of the rise to the run?

Answer by nerdybill(7384)

The ratio of the rise over the run is the "slope":
.
slope = rise/run
.
In your case, since
rise = 3/2
and
run = 5
.
then,
slope = rise/run = (3/2)/5 = 3/10